Calculate the molar solubility of Ag2SO4 when dissolved in water (Ksp = 1.9 x 10-5) (Please give your answer with 2 significant figures.)




1
Expert's answer
2021-09-15T02:41:34-0400



s="^3" √Ksp/4


"\sqrt[3]{\frac{Ksp}{4}}"

"\sqrt[3]{\frac{1.9×10^{-5}}{4}}"

= 0.01680987703

= 0.017